OpenStudy (anonymous):

What is the area of a circle with a radius of 2 meters? Use twenty-two sevenths for pi. three and one-seventh square meters six and two-sevenths square meters twelve and four-sevenths square meters twenty-five and one-seventh square meters

OpenStudy (anonymous):

@TylerD

OpenStudy (solomonzelman):

22/7 ≈ π A= πr²= (22/7)(2)²=(22/7)(4)=88/7=(12 × 7 + 4)/7=12 4/7